a conducting sphere of radius R, and carrying a charge q is joined to a conducting saphere of radius 2R and carrying a charge -2q. the charge flowing between them will be

Grade:11

1 Answers

Sher Mohammad IIT Delhi
9 years ago
the charge density density on two spheres will be same, q1/4pi*R^2=q2/4pi*(2R)^2, q2=4q1,
net charge =-q
5q1=-q, q1=-q/5, q2= -4q/5 , -q/5-q= -6q/5 , is the charge flow.


sher mohammad
b.tech, iit delhi